# SALAD: Achieve High-Sparsity Attention via Efficient Linear Attention Tuning for Video Diffusion Transformer

### v1

@@ 1 Introduction
Transformers face efficiency challenges in modeling long sequences due to the quadratic complexity of the attention mechanism [19]. Especially for the video diffusion transformer [20, 29], there is a challenge when generating high-resolution and long-duration videos. To reduce the inference cost of the video generation, various sparse attention mechanisms have been proposed [24, 28, 35, 40, 23, 37, 33]. Leveraging the inherent locality and sparsity of the attention map, these methods restrict each query to attend a subset of keys and values, thereby reducing the computation.
@@
Figure 1: Comparison of Our Method and Other Sparse Attention Mechanisms. Score versus speedup, with point size representing density—smaller points indicate lower computational density. The Summation Score is computed as summation of VBench metrics (Subject Consistency, Background Consistency, Imaging Quality, and Text Consistency). This reflects overall quality–efficiency trade-off. Models compared include our approach, SVG2 [28], PARO [40], ST-SWA [33], and ST-SWA + LoRA [7].
@@
Higher sparsity leads to lower computational cost in sparse attention. However, training-free sparse attention can only achieve limited sparsity [24, 28, 35, 33], such as $40\%\sim 60\%$ under sequence length of 30k. Training-based approaches can achieve remarkable sparsity of $80\%$ to $95\%$, yet they incur substantial training overhead in both data preparation and computation. For instance, VMoBA [23] requires approximately 182 GPU hours to train their sparse attention model on the large-scale Koala-36M dataset [22].
@@
Fine-tuning with Low-Rank Adaptation (LoRA) [7] provides a parameter-efficient approach for recovering performance after compression of large models [6, 36]. In Transformers, LoRA typically adapts the attention projection matrices ($W_{q}$, $W_{k}$, $W_{v}$, $W_{o}$) while keeping all other parameters frozen. But when applied to ultra-sparse attention models and fine-tuned with a limited training budget, achieving performance comparable to the dense counterpart becomes challenging. As shown in Fig. 1, the sparse attention model fine-tuned with LoRA still exhibits degradation relative to the full-attention baseline, even under a moderate sparsity level of 77.4%. The evaluation score is computed as the sum of selected VBench [8] metrics, including Subject Consistency, Background Consistency, Imaging Quality, and Text Consistency (Overall Consistency).
@@
Fig. 2 presents the results generated by the sparse attention model (specifically using a sliding-window mask), along with those from the full-attention and LoRA-tuned variants. While LoRA fine-tuning substantially improves the generative quality of the sparse attention model, the videos still exhibit artifacts inherently tied to the sparse pattern. For example, although the text prompt describes a dog, the LoRA-tuned model initially generates two dogs within the same frame, which gradually merge into one as the video progresses. Such textual and temporal inconsistencies highlight the limited effectiveness of tuning that relies solely on LoRA. Intuitively, these artifacts arise from the intrinsic limitations of sparse attention, which restricts cross-token interactions and leads to information loss. For example, sliding-window attention constrains each token to attend only to its local neighbors within a single layer. Even with multiple stacked layers, the receptive field cannot effectively expand to cover the entire sequence [25], thereby limiting the model’s ability to capture long-range dependencies. This incomplete context aggregation manifests as the text and temporal inconsistencies observed in Fig. 2.
@@
In principle, LoRA can compensate for the loss of cross-token interactions through its multi-layer stacked computations. During fine-tuning, token representations in each attention block are refined, enabling the current layer to incorporate information from distant tokens. Through layer-by-layer propagation, deeper layers can thus access broader contextual information, theoretically restoring long-range interactions. Consequently, LoRA fine-tuning holds the potential to recover performance comparable to full attention. However, our experiments show that achieving substantial recovery remains difficult under limited training budgets.
@@
However, we empirically find that although the linear attention branch can help supplement the lost information, it struggles to model the whole long sequences. Thus, the sparse attention should dominate the output while the linear attention branch should be set to an auxiliary role.
To achieve a practical sparse-linear attention , we propose SALAD (High-Sparsity Attention paralleling with Linear Attention for Diffusion Transformer), an efficient attention module to help achieve high-sparsity attention via linear attention branch. We first introduce a linear attention branch paralleling with the sparse attention. And then we design an Input-Dependent Scalar Gate to finely control the impact of linear attention layer by layer. Our contribution can be summarized as follows:
@@
    We observe that in hybrid sparse-linear attention, it is essential to regulate the influence of the linear attention branch. To this end, we design an input-dependent scalar gate that precisely controls its contribution, effectively improving overall model performance.
@@ 2 Related Work
2.1 Video Diffusion Models
Diffusion Transformer [17] has demonstrated its advantage in video generation. Earlier approaches such as Latte [15] decouple video sequences into spatial and temporal tokens and apply attention to each separately, which limits the model’s expressive capacity. More recent works [1, 29, 20] instead process the entire video as a unified token sequence and perform full attention across all tokens. While this strategy significantly improves generation quality, it also introduces substantial memory and computational overhead due to the extended sequence length.
2.2 Efficient Video Diffusion Transformer
To address the computational challenges in video diffusion transformers, various efficient video generation strategies have been proposed.
Quantization [39, 10, 3] reduces memory usage and latency in video diffusion models by lowering weight and activation precision. Caching mechanisms [14, 9, 5, 12] can reuse intermediate features across denoising timesteps, alleviating redundant computations. Timestep distillation [41] speeds up video generation by reducing the number of required sampling steps. Token merging compresses sequence length by combining similar tokens, mitigating memory and compute overheads for long videos [31].
Additionally, efficient attention methods have been explored a lot to reduce the computation caused by the long video token sequence. Sparse attention reduces the costs by limiting cross-token interactions [24, 32, 38, 28, 40]. Linear attention mechanism is also adopted. It transforms the quadratic complexity of attention to linear complexity, effectively reducing the burden of modeling the long video token sequence [21, 2].
@@ 2.2 Efficient Video Diffusion Transformer
2.3 LoRA for Model Compression Recovery
LoRA [7] is an efficient fine-tuning technique that freezes the pre-trained model weights and injects trainable low-rank decomposition matrices. Beyond downstream task adaptation, LoRA has also been widely adopted for model compression recovery. For instance, Ma et al. [16] combine LoRA with Large Language Models pruning to mitigate performance degradation with minimal fine-tuning overhead. QLoRA [6] introduces quantized LoRA tuning for 4-bit model performance recovery, while LongLoRA [4] extends LoRA to normalization, embedding and Feed-Forward Network (FFN) layers for more effective long-context adaptation under local attention. In this work, we find that LoRA fails to fully recover the performance of ultra-sparse (90%) Video Diffusion Transformers and propose SALAD as a more effective tuning approach.
@@ 3 Preliminary
3.1 Sparse Attention
Given queries ($Q$), keys ($K$), and values ($V$) $\in\mathbb{R}^{N\times d}$, the full attention mechanism can be formalized as follows:
@@ 3.1 Sparse Attention
$\mathrm{SparseAttention}(Q,K,V)=\operatorname{softmax}\left(\frac{QK^{\top}+M}{\sqrt{d}}\right)V.$ (2)
@@
Static sparse attention predefined the sparse mask prior to inference, often through heuristic or data-driven calibration strategies [32, 33, 40] that exploit prior knowledge about attention patterns. In contrast, dynamic approaches determine the sparse mask on-the-fly during inference, allowing the sparsity pattern to adapt to the input data via learned or rule-based decision functions [24, 28].
@@
Sliding-window attention (SWA) exploits the inductive bias of locality within the input sequence [32, 38]. Its sparse mask can be efficiently derived through calibration on a small subset of data. Motivated by [24], we further apply spatial-temporal reordering on this design to better handle video tokens, termed ST-SWA. Details of the ST-SWA used in our experiments are provided in the Appendix A.1.
Top-$K$ attention dynamically constructs irregular sparse masks by retaining the most informative query–key pairs. For each query, only the $k$ keys with the highest importance scores are preserved, allowing the model to focus computation on salient interactions. In this work, we use Top-K sparse attention implementation of VMoBA [23]. In the meantime, Top-$P$ attention adopts a similar principle but retains a variable number of keys whose cumulative importance exceeds a threshold $p$ [13, 30, 34].
3.2 Linear Attention
Linear attention reduces the computational complexity from $O(N^{2}d)$ in standard attention to $O(Nd^{2})$ by replacing the softmax operation with a kernel function $\phi(\cdot)$ applied to $Q$ and $K$. This allows the attention weights to be approximated as $\phi(Q)\phi(K)^{\top}$, enabling associative reordering. Specifically, one can first compute $H=\phi(K)^{\top}V$ and $Z=\phi(K)^{\top}\mathbf{1}$, and then obtain the output as $O=\frac{\phi(Q)H}{\phi(Q)Z}$.
A common choice for $\phi(\cdot)$ is ReLU function [26, 2]. The naive relu-based linear attention can be formulated as:
$O_{i}$ $=\sum_{j=1}^{N}\frac{\mathrm{ReLU}(Q_{i})\mathrm{ReLU}(K_{j})^{\top}V_{j}}{\sum_{j=1}^{N}\mathrm{ReLU}(Q_{i})\mathrm{ReLU}(K_{j})^{\top}}$
$=\frac{\mathrm{ReLU}(Q_{i})\left(\sum_{j=1}^{N}\mathrm{ReLU}(K_{j})^{\top}V_{j}\right)}{\mathrm{ReLU}(Q_{i})\left(\sum_{j=1}^{N}\mathrm{ReLU}(K_{j})^{\top}\right)}.$ (3)
@@ 4 Methods
4.1 Overview
The architecture of SALAD is presented in Fig.3. SALAD introduces a parallel linear attention branch to sparse attention, with the output of the linear attention added to the sparse attention output. The linear attention branch shares query (Q), key (K) and value (V) with sparse attention. The attention operation means the interaction of Q, K and V. The linear attention is followed by a projection layer. Following the prior design [34], we attempt to leverage it to balance the distribution of the two types of attention output.
SALAD can be generally applied to various sparse attention mechanisms, including both static and dynamic variants. In our experiments, we adopt a modified sliding-window attention [32] and a self-implemented top-$k$ sparse attention [23], as described in Sec. 3.1. These two approaches are the static and dynamic sparse attention, respectively. For the linear attention branch, we use a ReLU-based linear attention formulation.
For the linear attention branch, we observe that carefully regulating its output range is critical to overall performance. To achieve adaptive control over this branch, we introduce an input-dependent scalar gate. Specifically, the input hidden states are projected to a scalar that modulates the output of the linear attention before it is fused with the sparse attention branch. This design enables the model to dynamically adjust the contribution of the linear branch based on the input. The complete computation of the SALAD attention is presented in Algorithm 1.
Input: $X$: Input hidden states to the attention block
Output: $O$: Output of the attention block
$\mathbf{Q}=XW^{Q}$;
$\mathbf{K}=XW^{K}$;
$\mathbf{V}=XW^{V}$;
$O_{s}=\operatorname{SparseAttn}(\mathbf{Q},\mathbf{K},\mathbf{V})$ // Corresponds to Equation 2
$O_{l}=\operatorname{LinearAttn}(\mathbf{Q},\mathbf{K},\mathbf{V})$ // Corresponds to Equation 3
$G=\sigma(W_{G}X+b_{G})$ // $G\in\mathbb{R}^{n\times 1}$
$\bar{G}=\frac{1}{n}\sum_{i=1}^{n}G_{i}$;
$\tilde{O}=O_{s}+\bar{G}\cdot\operatorname{Proj}(O_{l})$
// $\operatorname{Proj}$ is a linear layer; $\bar{G}$ is the scalar GATE
@@ 4.1 Overview
// $W^{O}$ is a linear projection
Algorithm 1 SALAD Attention Computation
Next, Sec. 4.2 will provides an overview of the linear attention branch and the overall training pipeline of SALAD. Sec. 4.3 gives more details of GATE designing for precisely controlling the influence of linear attention.
4.2 Linear Attention Branch for Sparse Attention
For the linear attention branch, we adopt the naive ReLU-based linear attention formulation mentioned in Sec. 3.2. However, the original linear attention is not designed for video sequence modeling. Consequently, to handle the spatial-temporal relationship in video sequence, we integrate the 3D Rotary Position Embeddings (3D RoPE) [18] to linear attention branch. As Fig.3 presented, we apply the 3D RoPE directly to Query and Key, following what full attention dose [20].
SALAD introduces a few additional parameters. Specifically, a projection layer is used after the linear attention branch, following the design of [34]. The layer plays a role in balancing the distribution of the two types of attention outputs. And to obtain the gate, we add a linear layer to SALAD, which performs channel mixing on the features.
The weights $W_{q}$, $W_{k}$, $W_{v}$, and $W_{o}$ are initialized from the pretrained model and fine-tuned using LoRA. For the projection layer ($\operatorname{proj}$) in the linear attention branch, we apply zero-initialization to enable training to start from the vanilla sparse attention model. Both the projection layer weights and the input-dependent gate weights are fully trainable during fine-tuning.
Figure 4: The Rank of Sparse Attention and Linear Attention.
4.3 Input Dependent Scalar Gate
Prior studies [34] employ a single projection linear layer to integrate the outputs from the sparse and linear attention branches. Nevertheless, we empirically find that it is not sufficient to balance the output distribution of the branches.
Fig. 4 shows that the output rank of the linear attention branch is substantially lower than that of sparse attention, indicating a clear capacity imbalance. In the sparse–linear attention design, the sparse branch carries the primary sequence modeling load, while the linear branch functions as an auxiliary path. Since the linear branch mainly compensates for information missed by highly sparse attention, we speculate that properly regulating the contribution of the linear branch, rather than allowing it to dominate or vanish, is important for ensuring its effective complementarity to sparse attention.
Figure 5: Effect of Scaling the Linear Attention Branch. As $\lambda$ decreases, we observe improvements in background consistency, imaging quality, and text consistency. This suggests that the contribution of the linear attention branch to the final output must be carefully constrained. Relying solely on a static projection layer to regulate its influence appears insufficient, highlighting the need for more dynamic control over the branch fusion ratio.
However, we observe that introducing a hyperparameter $\lambda$ to modulate the linear attention branch can yield additional performance gains. This hyperparameter is incorporated as follows:
$O=O_{s}+\lambda\cdot\operatorname{Proj}(O_{l})$ (4)
As illustrated in Fig. 5, although $\lambda$ is set to $1.0$ during training, applying a moderately reduced scaling factor during inference improves performance in several aspects. But excessively small values of $\lambda$ will lead to performance degradation.
Intuitively, the role of the branch might vary across different layers and inputs. However, the projection layers hold a static weight, and therefore cannot adapt to different input conditions. This motivates us to design an Input-Dependent Scalar Gate for linear attention branch.
Fig.3 and Alg. 1 illustrate the gating operation.The input sequence of the branch is passed through a linear layer followed by a nonlinear activation. After that, we leverage token averaging to get a scalar gate and the linear attention output will be multiplied by the gate. Since the gate with scale smaller than one achieves a better performance, we adopt the sigmoid function as the non-linear function.

### v2

@@ 1 Introduction
Transformers face efficiency challenges in modeling long sequences due to the quadratic complexity of the attention [vaswani2017attention]. Especially for the video diffusion transformer [wan2025, yang2024cogvideox], there is a challenge when generating high-resolution and long-duration videos. To reduce the cost of video generation, various sparse attention mechanisms have been proposed [xi2025sparse, yang2025sparse, zhang2025spargeattn, zhao2025paroattention, wu2025vmoba, zhang2025vsa, zhang2025ditfastattnv2]. Leveraging the inherent locality and sparsity of the attention map, these methods restrict each query to attend a subset of keys and values, thereby reducing the computation.
@@
Figure 1: Comparison of SALAD and Other Sparse Attention Mechanisms on Wan1.3B [wan2025] with 30k tokens length. VBench Score versus speedup, with point size representing density—smaller points indicate lower computational density. This reflects overall quality–efficiency trade-off. Models compared include our approach, SVG2 [yang2025sparse], PARO [zhao2025paroattention], and ST-SWA [zhang2025ditfastattnv2] + LoRA [hu2022lora].
@@
Higher sparsity leads to lower computational cost in sparse attention. However, training-free sparse attention can only achieve limited sparsity [xi2025sparse, yang2025sparse, zhang2025spargeattn, zhang2025ditfastattnv2], such as $40\%\sim 60\%$ under sequence length of 30k. Training-based approaches can achieve remarkable sparsity of $80\%$ to $95\%$, yet they incur substantial training overhead in both data preparation and computation. For instance, VMoBA [wu2025vmoba] requires approximately 182 GPU hours to train their sparse attention model on the large-scale Koala-36M dataset [wang2025koala].
@@
Fine-tuning with Low-Rank Adaptation (LoRA) [hu2022lora] provides a parameter-efficient approach for recovering performance after compression of large models [dettmers2023qlora, zhang2023loraprune]. In Transformers, LoRA typically adapts the attention projection matrices ($W^{Q}$, $W^{K}$, $W^{V}$, $W^{O}$) while keeping all other parameters frozen. But when applied to ultra-sparse attention models and fine-tuned with a limited training resources, achieving performance comparable to the dense counterpart becomes challenging. As shown in Fig. 1, the sparse attention model fine-tuned with LoRA still exhibits degradation relative to the full-attention baseline, even under a moderate sparsity. The evaluation score is computed as the sum of selected VBench [huang2024vbench] metrics, including Subject Consistency, Background Consistency, Imaging Quality, and Text Consistency (Overall Consistency).
@@
Fig. 2 presents the results generated by the sparse attention model (specifically using a sliding-window mask), along with those from the full-attention and LoRA-tuned variants. While LoRA fine-tuning substantially improves the generative quality of the sparse attention model, the videos still exhibit artifacts inherently tied to the sparse pattern. For example, although the text prompt describes a dog, the LoRA-tuned model initially generates two dogs within the same frame, which gradually merge into one as the video progresses. Such textual and temporal inconsistencies highlight the limited effectiveness of tuning sparse attention relies solely on LoRA. Intuitively, these artifacts arise from the intrinsic limitations of sparse attention, which restricts cross-token interactions and leads to information loss. For example, sliding-window attention constrains each token to attend only to its local neighbors within a single layer. Even with multiple stacked layers, the receptive field cannot effectively expand to cover the entire sequence [xiao2025sliding], thereby limiting the model’s ability to capture long-range dependencies. This incomplete context aggregation manifests as the text and temporal inconsistencies observed in Fig. 2.
@@
In principle, LoRA can compensate for the loss of cross-token interactions through multi-layer stacked computations. During fine-tuning, token representations in each attention block are refined, enabling the current layer to incorporate information from distant tokens. Through layer-by-layer propagation, deeper layers can thus access broader contextual information, theoretically restoring long-range interactions. Consequently, LoRA fine-tuning holds the potential to recover performance comparable to full attention. However, our experiments show that achieving substantial recovery remains difficult under extreme high sparsity.
@@
However, we empirically find that the linear attention branch can only supplement limited critical information while struggling to model the entire sequence. Thus, the sparse attention should dominate the output while the linear attention branch should be set to an auxiliary role.
To enable practical sparse-linear attention, we propose SALAD (High-Sparsity Attention paralleling with Linear Attention for Diffusion Transformer), an efficient attention module that facilitates high-sparsity modeling through an auxiliary linear attention branch. Specifically, we introduce a linear attention branch in parallel with sparse attention. We further design a Multi-level Static-Dynamic Scaling Strategy to regulate the linear branch across different blocks and timesteps in the Diffusion Transformer. Our contribution can be summarized as follows:
@@
    We observe that in hybrid sparse-linear attention, it is essential to regulate the influence of the linear attention branch. We design a Multi-level Static-Dynamic Scaling Strategy that precisely controls its contribution, effectively improving overall model performance.
@@ 2 Related Work
Video Diffusion Models. Diffusion Transformer [Peebles2022DiT] has demonstrated its advantage in video generation. Earlier approaches such as Latte [ma2024latte] decouple video sequences into spatial and temporal tokens and apply attention separately, which limits the model’s expressive capacity. More recent works [brooks2024video, yang2024cogvideox, wan2025] instead process the entire video as a unified token sequence and perform full attention across all tokens. While improving generation quality, it introduces substantial memory and computational overhead due to the extended sequence length.
Efficient Video Diffusion Transformer. To address the computational challenges of video diffusion transformers, various efficient video generation techniques have been proposed. Quantization [zhao2024viditq, li2024svdquant, chen2025q] reduces memory usage and inference latency by lowering the precision of model weights and activations. Caching mechanisms [lv2024fastercache, kahatapitiya2025adaptive, cui2025bwcache, liu2025timestep] reuse intermediate features across denoising timesteps to avoid redundant computations. Timestep distillation [zheng2025large] accelerates generation by reducing the number of required sampling steps. Token merging compresses the sequence length by merging similar tokens, thereby mitigating the memory and computational overhead of long videos [yuan2025dlfr]. In addition, efficient attention mechanisms have been widely explored to address the long video token sequence. Sparse attention reduces computational cost by restricting cross-token interactions [xi2025sparse, yuan2024ditfastattn, zhang2025fast, yang2025sparse, zhao2025paroattention]. Linear attention has also been adopted, transforming the quadratic complexity of attention into linear complexity and thus reducing the cost of modeling long video sequences [wang2025lingen, chen2025sana].
@@
LoRA for Model Compression Recovery. LoRA [hu2022lora] is an efficient fine-tuning technique that freezes the pre-trained model weights and injects trainable low-rank decomposition matrices. Beyond downstream task adaptation, LoRA has also been widely adopted for model compression recovery. For instance, Ma et al. [ma2023llm] combine LoRA with Large Language Models pruning to mitigate performance degradation with minimal fine-tuning overhead. QLoRA [dettmers2023qlora] introduces quantized LoRA tuning for 4-bit model performance recovery, while LongLoRA [chen2023longlora] extends LoRA to normalization, embedding, and Feed-Forward Network layers for long-context adaptation under local attention. In this work, we find that LoRA fails to fully recover the performance of ultra-sparse Video Diffusion Transformers and propose SALAD as a more effective tuning approach.
@@ 3 Preliminary
Sparse Attention. Given queries ($Q$), keys ($K$), and values ($V$) $\in\mathbb{R}^{N\times d}$, the full attention mechanism can be formalized as follows:
@@
$\mathrm{SparseAttention}(Q,K,V)=\operatorname{softmax}\left(\frac{QK^{\top}+M}{\sqrt{d_{k}}}\right)V.$ (2)
@@
Static sparse attention predefined the sparse mask prior to inference, often through heuristic or data-driven calibration strategies [yuan2024ditfastattn, zhang2025ditfastattnv2, zhao2025paroattention] that exploit prior knowledge about attention patterns. In contrast, dynamic approaches determine the sparse mask on-the-fly during inference, allowing the sparsity pattern to adapt to the input data via learned or rule-based decision functions [xi2025sparse, yang2025sparse].
@@
Sliding-window attention (SWA) exploits the inductive bias of locality within the input sequence [yuan2024ditfastattn, zhang2025fast]. Its sparse mask can be efficiently derived through calibration on a small subset of data. Motivated by  [xi2025sparse], we further apply spatial-temporal reordering on this design to better handle video tokens, termed ST-SWA. Details of it are provided in the Appendix 0.A.1.
Top-$K$ attention dynamically constructs irregular sparse masks by retaining the most informative query–key pairs. For each query, only the $k$ keys with the highest importance scores are preserved, allowing the model to focus computation on salient interactions. In this work, we use Top-K sparse attention implementation of VMoBA [wu2025vmoba]. In the meantime, Top-$P$ attention adopts a similar principle but retains a variable number of keys whose cumulative importance exceeds a threshold $p$ [lu2025moba, yuan2025native, zhang2025sla].
Linear Attention. Linear attention reduces the computational complexity from $O(N^{2}d)$ in standard attention to $O(Nd^{2})$ by replacing the softmax operation with a kernel function $\phi(\cdot)$ applied to $Q$ and $K$. This allows the attention weights to be approximated as $\phi(Q)\phi(K)^{\top}$, enabling associative reordering. Specifically, one can first compute $H=\phi(K)^{\top}V$ and $Z=\phi(K)^{\top}\mathbf{1}$, and then obtain the output as $O=\frac{\phi(Q)H}{\phi(Q)Z}$. A common choice for $\phi(\cdot)$ is ReLU function [xie2024sana, chen2025sana]. The relu-based linear attention can be formulated as:
$O_{i}$ $=\frac{\mathrm{ReLU}(Q_{i})\left(\sum_{j=1}^{N}\mathrm{ReLU}(K_{j})^{\top}V_{j}\right)}{\mathrm{ReLU}(Q_{i})\left(\sum_{j=1}^{N}\mathrm{ReLU}(K_{j})^{\top}\right)}.$ (3)
@@ 4 Methods
The architecture of SALAD is illustrated in Fig. 3. To achieve high-sparsity attention in video diffusion transformers while preserving generation quality, SALAD augments sparse attention with a lightweight parallel linear attention branch. Both branches share the same query ($\mathbf{Q}$), key ($\mathbf{K}$), and value ($\mathbf{V}$). The projection matrices $W^{Q}$, $W^{K}$, $W^{V}$, and $W^{O}$ are initialized from the pretrained weights and fine-tuned using LoRA during training. The linear attention output is residually added to the sparse output after multi-leval scaling.
In this work, we observe that carefully regulating the output range of linear attention is critical to overall performance. Thus, we introduce a multi-level static-dynamic scaling strategy that finely balances the linear branch’s contribution. This comprises (1) a layer-wise static scaling implemented via a zero-initialized projection layer and (2) a lightweight input- and timestep-aware dynamic scaling module that modulates the linear influence based on the denoising timestep $t$ and current hidden states. Together, these mechanisms ensure the linear branch acts as a controlled complement, efficiently injecting missed information under extreme sparsity without overwhelming sparse attention output. The complete SALAD attention computation is presented in Alg. 1.
SALAD can be generally applied to various sparse attention mechanisms, including both static and dynamic variants. In our experiments, we adopt a modified sliding-window attention [yuan2024ditfastattn] and a self-implemented top-$k$ sparse attention [wu2025vmoba], as described in Sec. 3. These two approaches are the static and dynamic sparse attention, respectively. For the linear attention branch, we use a ReLU-based linear attention. The linear attention branch incorporates 3D Rotary Position Embeddings [su2024roformer] to model spatial-temporal dependencies.
Input: $X$: Unnormalized hidden states; $t$: Denoising timestep
Output: $O$: Output of the SALAD attention block
$e_{t}=\operatorname{TimeEmb}(t)$ // Time embedding inherited from the pretrained DiT
$X_{t}=\operatorname{Modulation}(\operatorname{norm}(X),e_{t})$
// Feature Transformation
$\mathbf{Q}=X_{t}W^{Q},\quad\mathbf{K}=X_{t}W^{K},\quad\mathbf{V}=X_{t}W^{V}$;
$O_{s}=\operatorname{SparseAttn}(\mathbf{Q},\mathbf{K},\mathbf{V})$
$O_{l}=\operatorname{LinearAttn}(\mathbf{Q},\mathbf{K},\mathbf{V})$
// Level 1: Static Layer-wise Scaling
$O^{\prime}_{l}=\operatorname{Proj}(O_{l})$ ;
// Zero-initialized linear layer
// Level 2: Dynamic Input-Timestep-Aware Scaling
$s=\operatorname{Mean}(\sigma(\operatorname{Linear}(X_{t}))$ ;
// Global scaling factor
// Residual Integration and Final Projection
$\tilde{O}=O_{s}+s\cdot O^{\prime}_{l}$
@@
Algorithm 1 SALAD Attention with Multi-level Scaling
4.1 Rank Imbalance between Linear and Sparse Branches
A straightforward way to enhance sparse attention with linear attention is to directly combine their outputs by summation and fine-tune the model. However, we empirically observe that this integration leads to suboptimal performance.
we analyze the attention maps and effective rank of the two branches. As shown in Fig. 5, the linear attention branch indeed captures global token interactions, suggesting its potential to compensate for information missed by highly sparse attention. Nevertheless, examining the effective rank of the branch outputs uncovers a significant imbalance. As illustrated in Fig. 6 (Left), the linear attention output exhibits a substantially lower rank than that of sparse attention, indicating a pronounced capacity gap between the two branches.
More importantly, we find that directly adding the two branches can even reduce the overall output rank compared to its of sparse attention. This suggests that the low-rank linear branch may inadvertently suppress informative high-rank structures from the sparse attention when its magnitude is not properly controlled. These observations indicate that the linear attn branch should serve as a auxiliary complement rather than a co-equal contributor. If over-amplified, it may overwhelm sparse attention output and degrade the modeled information.
Figure 4: Linear and sparse attention maps at block 8, timestep 20, head 2. The linear attention can achieve global token interaction, whereas sparse attention restricts each query to interact with only a subset of tokens.
Figure 5: Generated Video Samples of sparse attention tuned with LoRA, SALAD, and SALAD with linear branch dropped during inference.
Figure 6: Effective rank comparison and effect of linear attention scaling. (Left) Effective rank of sparse attention and linear attention outputs. (Right) As $\lambda$ decreases, we observe improvements in some dimensions. Moreover, the optimal value of $\lambda$ varies across different inputs. These observations suggest that relying solely on a static projection layer to regulate the influence of the linear attention branch is insufficient.
4.2 Multi-level Static-Dynamic Scaling for Linear Attention Branch
The rank imbalance analysis indicates that the sparse branch carries the primary sequence modeling load, while the linear branch functions as an auxiliary path, mainly compensating for information missed by highly sparse attention. We speculate that the low-rank linear attention component may suppress or overwhelm informative high-rank structures from the sparse branch without explicit regulation. Therefore, properly down-weighting the linear branch, rather than allowing it to dominate or vanish is crucial. To this end, we introduce a multi-level static-dynamic scaling strategy, consisting of a layer-wise static scaler and an input-timestep-aware dynamic scaler.
Layer-wise Static Scaling (LWSS). As shown in Fig. 6, the degree of rank degradation varies across different blocks. And different transformer layers exhibit varying sensitivity to sparse attention [yuan2024ditfastattn, zhang2025ditfastattnv2], and consequently require different degrees of compensation from the linear attention branch. To accommodate this heterogeneity, we introduce a layer-wise static scaling mechanism that adaptively down-weights the contribution of linear attention at each layer.
Experiments reveal that inserting a learnable projection layer after the linear attention branch in each Transformer layer helps preliminarily regulate the contribution of linear attention. And inspired by the LoRA [hu2022lora], the projection layer is zero-initialized. It allows the model to start tuning from the post-training sparse attention, ensuring more stable tuning in the early stage. Meanwhile, this design encourages the linear attention branch to be progressively incorporated from a small-magnitude regime, enabling and gradual increase of its contribution. The projection layer is trained during fine-tuning and is formulated as:
$O=O_{s}+\operatorname{ZeroInitProj}(O_{l}),$ (4)
where $O_{s}$ and $O_{l}$ denote the outputs of the sparse and linear attn, respectively.
However, when introducing a scaling hyperparameter $\lambda$,
$O=O_{s}+\lambda\cdot\operatorname{ZeroInitProj}(O_{l}),$ (5)
we find that the performance can be further improved. As shown in Fig. 6 (Right), although models are trained with $\lambda=1$, moderately reducing $\lambda$ during inference often yields better results, while excessively small values degrade performance. This phenomenon suggests that the linear branch indeed requires more careful magnitude control. Relying solely on a static scaling strategy appears insufficient, highlighting the need for more dynamic control. However, both the projection layer and the hyperparameter $\lambda$ are static. They cannot adapt to different inputs or diffusion timesteps, resulting in a suboptimal performance.
Input-Timestep-Aware Dynamic Scaling (ITADS). To enable adaptive regulation, we introduce an input-timestep-aware dynamic scaling module. As depicted in Alg. 1 and Fig. 3, the input fused with the timestep embedding is passed to a learnable linear layer with channel mixing, followed by a nonlinear function. Then token averaging is performed to produce a global scaling factor.
Since scaling factors $s$ smaller than one empirically lead to better behavior, we adopt a sigmoid activation to constrain $s\in(0,1)$. The dynamic scaling module introduces minimal overhead and is fully trainable during fine-tuning.
Together, the static and dynamic scaling mechanisms ensure that the linear branch remains an auxiliary low-rank complement, while adaptively injecting global information when necessary. After applying the proposed scaling strategy, we observe that the rank of the combined output remains comparable to sparse attention output. Moreover, we observe a lower subspace overlap between the linear and sparse branches, measured by the Frobenius norm of the projection between their top principal components obtained via PCA [mackiewicz1993principal]. This suggests that the two branches learn more complementary representations, enabling the linear branch to effectively compensate for information missing in sparse attention. More details are provided in the Appendix 0.A.2.
4.3 Post-tuning Branch Dropping
The contribution of the linear attention branch varies across timesteps and inputs, as reflected by the dynamic scaling factors. We observe that branches with small scaling values contribute marginally and can be dropped without degrading performance. Based on this observation, we introduce a simple branch-dropping strategy guided by the scaling factors.
Specifically, we calibrate on a small validation prompt set to determine a timestep-wise threshold $\tau_{t}$. During inference, branches with global scaling factor below $\tau_{t}$ are dynamically pruned. More details are provided in the Appendix 0.A.3.
Interestingly, dropping a small portion of branches with extremely low scaling values can even slightly improve generation quality (see Table 2). We conjecture that some layers at a certain timestep are naturally better suited for sparse attention, such that the linear branch may introduce minor negative effects.